OpenStudy (anonymous):

If a stone is tossed from the top of a 190 meter building, the height of the stone as a function of time is given by h(t) = -9.8t2 – 10t + 190, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer.

OpenStudy (radar):

geekgirl1988, please review the previous question you asked regarding the application of the discriminant. I made an error which I corrected.

OpenStudy (anonymous):

-9.8t^2 - 10t + 190 = 190 t(-9.8t-10) = 0 t=-10/9.8 Thus it will fall after 10/9.8 secs

OpenStudy (anonymous):

yes i saw that radar and it is one solution.

OpenStudy (anonymous):

I think i did a mistake...sorry!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The equation you need to solve is -9.8t^2 - 10t + 190 = 0

OpenStudy (anonymous):

would i use the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok on while i solve the problem.

OpenStudy (anonymous):

you know the formula right

OpenStudy (anonymous):

yes. doing th problem right now. almost done. hold on.

OpenStudy (anonymous):

After you do it, just tell me how old are you? So that I can help accordingly, or else i might miss something you may not know

OpenStudy (anonymous):

i got x=-3.6 or x=2.6 for my answer. and i am 22.

OpenStudy (anonymous):

Right, now the point is can time be negative?

OpenStudy (anonymous):

would that answer be no. if so then the answer to the problem would be 2.6 seconds per meter so it would be 494 seconds. or am i wrong.

OpenStudy (anonymous):

Yup you are right, btw how did you get that value 494?

OpenStudy (anonymous):

2.6*190=494. should i not put that into the answer.

OpenStudy (anonymous):

Nope, that doesn't mean anything!

OpenStudy (anonymous):

The last part would be 2.667